# Out-of-Distribution Semantic Occupancy Prediction

### v1

@@ Abstract
3D Semantic Occupancy Prediction is crucial for autonomous driving, providing a dense, semantically rich environmental representation. However, existing methods focus on in-distribution scenes, making them susceptible to Out-of-Distribution (OoD) objects and long-tail distributions, which increases the risk of undetected anomalies and misinterpretations, posing safety hazards. To address these challenges, we introduce Out-of-Distribution Semantic Occupancy Prediction, targeting OoD detection in 3D voxel space. To fill the gaps in the dataset, we propose a Synthetic Anomaly Integration Pipeline that injects synthetic anomalies while preserving realistic spatial and occlusion patterns, enabling the creation of two datasets: VAA-KITTI and VAA-KITTI-360. We introduce OccOoD, a novel framework integrating OoD detection into 3D semantic occupancy prediction, with Voxel-BEV Progressive Fusion (VBPF) leveraging an RWKV-based branch to enhance OoD detection via geometry-semantic fusion. Experimental results demonstrate that OccOoD achieves state-of-the-art OoD detection with an AuROC of 67.34% and an AuPRCr of 29.21% within a 1.2m region, while maintaining competitive occupancy prediction performance. The established datasets and source code will be made publicly available at https://github.com/7uHeng/OccOoD.
1 Introduction
@@
3D scene understanding is fundamental to autonomous driving, crucial for motion planning and obstacle avoidance [36]. The cost-effectiveness and ease of deployment of cameras have made vision-based 3D understanding methods a focal point of research [12, 66, 121, 37, 114, 97, 82, 53, 62]. Among these methods, 3D semantic occupancy prediction networks [41, 51, 66] are distinguished by their ability to accurately reconstruct geometry and semantic details. By producing detailed occupancy maps with rich semantic context, these networks improve environmental perception and decision-making, enhancing the safety and reliability of autonomous navigation.
@@
However, current methods [41, 85, 51] primarily focus on in-distribution scenarios, where models are trained and evaluated on predefined object categories and environments. While effective in controlled settings, this approach [53, 12] exposes systems to vulnerabilities from OoD objects and anomalies commonly encountered in real-world scenarios [55, 8, 33, 17], as seen in Fig. LABEL:fig:teaser where the standard semantic occupancy prediction method [66] produces misclassifications. OoD detection—the capability to identify entities that substantially deviate from the training data distribution—remains a critical yet underexplored challenge. Such anomalies include unexpected obstacles (e.g., discarded furniture, stray wildlife, construction equipment), unusual road conditions (e.g., potholes, debris, or oil spills), and atypical weather phenomena (e.g., sudden fog, heavy rain, or snowdrifts). As existing systems may misinterpret or fail to detect these deviations, they pose significant safety risks, potentially leading to failures and even catastrophic consequences in autonomous driving.
@@
To bridge this critical gap, we introduce the task of Out-of-Distribution Semantic Occupancy Prediction. However, a key challenge in this domain is the scarcity of datasets annotated for both semantic occupancy and OoD anomalies, hindering robust model development. To address this, we propose a Synthetic Anomaly Integration Pipeline that injects diverse anomalies into existing datasets while preserving realistic spatial and occlusion patterns. This pipeline enables the creation of two novel datasets comprising $26$ distinct categories of anomalies, where anomaly annotations encompass a diverse range of unforeseen hazards, facilitating comprehensive evaluation of OoD detection in 3D scene understanding. Building upon our newly proposed datasets for out-of-distribution detection in semantic occupancy prediction, we introduce OccOoD, the first unified framework that seamlessly integrates OoD detection into 3D semantic occupancy prediction.
@@
OccOoD incorporates several key innovations to enhance anomaly detection in complex environments. First, existing methods [66, 51] rely on a single-voxel representation, which limits the effectiveness of OoD detection. To address this, we propose a Voxel-BEV Progressive Fusion strategy, which integrates voxel-based and Bird’s-Eye-View (BEV) representations. BEV enhances global scene understanding, enabling improved occupancy prediction and robust OoD detection. Second, we introduce a dedicated OoD detection mechanism leveraging entropy and cosine similarity, enabling effective identification of anomalies in dynamic and unstructured scenes. Finally, to ensure comprehensive evaluation, we conduct extensive OoD detection experiments on the newly proposed datasets and verify our occupancy prediction performance based on public benchmarks. The experimental results show that OccOoD achieves AuROC scores of $65.91\%$ and $67.34\%$ on the two proposed datasets, respectively, and attains mIoU scores of $15.85\%$ and $18.37\%$ on SemanticKITTI [3], and KITTI-360 [52], respectively. These results collectively establish OccOoD as a robust and scalable solution for addressing OoD challenges in 3D semantic occupancy prediction.
In a nutshell, our contributions are three-fold:
@@
  • We for the first time introduce out-of-distribution detection into the task of semantic occupancy prediction by delivering two new datasets VAA-KITTI and VAA-KITTI-360, specifically designed to evaluate out-of-distribution detection in the context of autonomous driving safety.
@@
  • We propose OccOoD, a multi-view fusion framework that integrates front-view and bird’s-eye-view representations, enabling a comprehensive spatial understanding of outliers and seamless integration of OoD detection into 3D semantic occupancy prediction. This framework robustly handles both known and unknown objects.
@@
  • Our proposed framework, OccOoD, achieves state-of-the-art OoD detection performance on the established datasets while maintaining competitive accuracy in 3D semantic occupancy prediction.
2 Related Work
@@
Semantic Occupancy Prediction. 3D Semantic Occupancy Prediction aims to infer both the geometry and semantics of a scene by inferring the occupancy state of voxels [100, 73, 19, 79, 105]. To advance estimation quality and efficiency, efforts are made in 3D representation [23, 56, 62, 35, 78, 98, 82, 117, 39, 49, 61, 44, 42], context enhancement [101, 86, 46, 89, 124, 96, 48], modality fusion [30, 113, 70, 14, 75, 64, 104, 120], temporal association [92, 69, 91, 109, 108, 11], long-sequence and generative modeling [54, 90, 123, 47, 85], self-supervision [38, 32, 10, 9, 71], and offboard processing [63, 77, 94, 118, 40]. These efforts have collectively improved the quality and efficiency of semantic occupancy prediction. Recently, several works [121, 41, 66, 110] have introduced more fine-grained feature processing techniques. Meanwhile, uncertainty-aware methods [93, 80, 13] incorporate uncertainty quantification to enhance estimation reliability. However, these methods are limited to predefined categories, requiring retraining to generalize to novel object types and lacking flexibility. While open-vocabulary learning [81, 119, 9, 87, 116] addresses generalization to arbitrary categories, it remains constrained to In-Distribution (ID) scenarios and fails to handle OoD detection. In real-world applications, the categories of anomalous objects are unforeseeable, making such methods less reliable. In contrast, our proposed model extends semantic occupancy prediction by perceiving all OoD objects, enabling efficient and comprehensive OoD detection.
@@
Figure 1: An illustration of the proposed Synthetic Anomaly Integration Pipeline. This pipeline is designed to address the challenges of collecting real-world OoD data by synthesizing anomalies that are both physically plausible and contextually realistic.
@@
Out-of-Distribution Segmentation. Out-of-distribution segmentation identifies anomalies in specific scenes [22, 17, 55, 8] while maintaining in-distribution segmentation capabilities. Current research primarily focuses on anomaly detection in 2D images of road scenes, with the main methods including per-pixel architectures and mask-transformer-based approaches [1, 29, 74, 68, 25]. The per-pixel architecture methods include uncertainty-based [67, 21, 33, 43, 16], reconstruction-based [55, 99, 31, 88, 27], and outlier exposure [18, 84, 28, 58, 20] methods. Although these methods have advanced 2D anomaly detection, they are limited to narrow field-of-view pinhole cameras, hindering their application to complex 3D scenarios requiring high-precision spatial perception. Some studies extend anomaly detection to 3D space; however, these efforts primarily focus on industrial [5, 57] or indoor object defect detection [7, 50], where models fuse static point cloud data to optimize defect localization [34, 4, 15, 122, 111]. Due to reliance on geometric consistency assumption of static objects and limited sensitivity to dynamic changes, these methods fail to meet the demands of more complex scenarios. In comparison, our work introduces OoD segmentation into vision-based 3D semantic occupancy prediction, overcoming the limitations of current methods in dynamic environments.
3 Out-of-Distribution Datasets
We introduce the task of Out-of-Distribution Semantic Occupancy Prediction. To address the lack of evaluation platforms, we propose a Synthetic Anomaly Integration Pipeline to generate abnormal objects in real road scene images, leading to the creation of two novel datasets: VAA-KITTI and VAA-KITTI-360.
3.1 Synthetic Anomaly Integration Pipeline
@@
Collecting images and point cloud data of OoD objects is challenging due to their rarity in real-world scenes and the high cost of data collection and annotation. To address this, we propose a Synthetic Anomaly Integration Pipeline that generates synthetic anomalies under physical and environmental constraints, ensuring plausibility and challenge for robust OoD detection model evaluation. As illustrated in Fig. 1, our pipeline is structured into three key phases:
@@
Phase 1: Anomaly Generation and Localization. In 2D anomaly detection, studies have demonstrated that synthetic anomaly data is highly effective for model training and evaluation [24, 59]. We employ POC [24], a state-of-the-art image generation technique, to synthesize anomaly images and their corresponding mask maps, enabling precise localization of anomalous objects. To better align with real-world application scenarios, we introduce multi-category anomalies into the images, as shown in Fig. 2.
@@
Phase 2: Depth Estimation and 3D Transformation. This phase focuses on integrating image data with accurate depth information, facilitating the transformation of 2D anomaly detection results into precise 3D spatial localization. To address the challenge of limited metric depth estimation accuracy, we propose a method to project synthesized anomalies into the 3D point cloud.
@@
We first utilize the large version of Depth Anything V2 [103], a state-of-the-art depth estimation model, to compute pseudo-depth for both original and synthesized images. Subsequently, we randomly sample a predefined number of points, from the point cloud and project them onto the image plane to obtain pixel coordinates $\mathbf{u}_{i}$ and Euclidean distances $d_{i}$ as:
@@
where 00 and $1$ represent the occupancy status of the voxel.
3.2 VAA-KITTI and VAA-KITTI-360
Figure 2: Distribution map of VAA-KITTI and VAA-KITTI-360 datasets. The datasets contain $26$ distinct OoD categories, which can be divided into five main categories as shown in the figure.
We applied the aforementioned synthesis pipeline to the SemanticKITTI [3] and SSCBench-KITTI-360 [52] datasets, resulting in the creation of two new synthetic datasets: VAA-KITTI and VAA-KITTI-360. These synthesized datasets extend the original label sets by introducing anomaly labels, encompassing $26$ distinct categories of anomalies such as animals, furniture, and garbage bags.
To ensure the plausibility and naturalness of the generated images, over $30,000$ images produced using the POC algorithm were carefully reviewed by three independent annotators. After multiple rounds of cross-validation, images with poor generation quality, implausible object placements, or non-anomalous scenarios (e.g., birds flying in an urban sky or trash, trash bags, and wheeled bins placed alongside sidewalks) are excluded, resulting in a final selection of $500$ images for both the VAA-KITTI and VAA-KITTI-360 datasets to form the test set for the Out-of-Distribution Semantic Occupancy Prediction task.
Regional-Level Metrics Design: Referring to the field of 2D image anomaly detection, we adopt Area under the Receiver Operating Characteristic Curve (AuROC) as one of the evaluation metrics for model performance. Furthermore, to mitigate the evaluation impact caused by the inaccuracy of existing methods in predicting the absolute positions of instances, and bridge the gap between localization demands and anomaly-awareness goals, we propose a regional-level evaluation metric: Regional Area under the Precision-Recall Curve ($\text{AuPRC}_{r}$). The $\text{AuPRC}_{r}$ incorporates tolerance for minor localization inaccuracies while maintaining sensitivity to anomaly detection.
$\text{AuPRC}_{r}$: The Precision-Recall (PR) curve evaluates the balance between precision and recall at varying thresholds:
@@ 3.2 VAA-KITTI and VAA-KITTI-360
4 OccOoD: Proposed Framework
4.1 Problem Setting
@@
Out-of-distribution segmentation in 2D pinhole images typically relies on pixel-wise classification scores. Building on this concept, we extend out-of-distribution detection to 3D semantic occupancy prediction at the voxel level. Given an input image $I\in\mathbb{R}^{C\times H\times W}$, the semantic occupancy prediction process yields a volumetric representation $V\in\mathbb{R}^{C\times H\times W\times Z}$, where $Z$ denotes the depth dimension. The voxel-level classification scores $Q(V)\in\mathbb{R}^{K\times H\times W\times Z}$ are then obtained. To assess anomalies at the voxel level, an anomaly score $A(V)\in\mathbb{R}^{H\times W\times Z}$ is computed as Eq. (6):
$A(V)=f(Q(V)),$ (6)
where $K$ denotes the number of classes for the semantic prediction, $C$, $H$, $W$, and $Z$ correspond to the channels, height, width, and depth of the voxel space, respectively.
4.2 Overall Architecture

### v2

@@ Abstract
3D semantic occupancy prediction is crucial for autonomous driving, providing a dense, semantically rich environmental representation. However, existing methods focus on in-distribution scenes, making them susceptible to Out-of-Distribution (OoD) objects and long-tail distributions, which increases the risk of undetected anomalies and misinterpretations, posing safety hazards. To address these challenges, we introduce Out-of-Distribution Semantic Occupancy Prediction, targeting OoD detection in 3D voxel space. To fill dataset gaps, we propose a Realistic Anomaly Augmentation that injects synthetic anomalies while preserving realistic spatial and occlusion patterns, enabling the creation of two datasets: VAA-KITTI and VAA-KITTI-360. Then, a novel framework that integrates OoD detection into 3D semantic occupancy prediction, OccOoD, is proposed, which uses Cross-Space Semantic Refinement (CSSR) to refine semantic predictions from complementary voxel and BEV representations, improving OoD detection. Experimental results demonstrate that OccOoD achieves state-of-the-art OoD detection with an AuROC of $65.50\%$ and an AuPRCr of $31.83\%$ within a $1.2m$ region, while maintaining competitive semantic occupancy prediction performance and generalization in real-world urban driving scenes. The established datasets and source code will be made publicly available at https://github.com/7uHeng/OccOoD.
I Introduction
@@
3D scene understanding is fundamental to autonomous driving, crucial for motion planning and obstacle avoidance [18]. The cost-effectiveness and ease of deployment of cameras have made vision-based 3D understanding methods a focal point of research [9, 53]. Among these methods, 3D semantic occupancy prediction networks [26] are distinguished by their ability to accurately reconstruct geometry and semantic details. By producing detailed occupancy maps with rich semantic context, these networks improve environmental perception and decision-making, enhancing the safety and reliability of autonomous navigation.
@@
However, current methods [50, 35, 22, 40, 34, 32] primarily focus on in-distribution scenarios, where models are trained and evaluated on predefined semantic distributions and scenes. While effective in controlled settings, this approach [28] exposes systems to vulnerabilities from OoD objects and anomalies commonly encountered in real-world scenarios [7, 10], as seen in Fig. 1 where the standard semantic occupancy prediction method [36] produces misclassifications. OoD detection, the ability to identify entities that substantially deviate from the training data distribution, remains a critical yet underexplored challenge.
Figure 1: Visualization of the established task of Out-of-Distribution Semantic Occupancy Prediction. The mainstream semantic occupancy prediction methods tend to misclassify Out-of-Distribution (OoD) objects as inliers, thus endangering the safety of autonomous driving. Proposed OccOoD accurately identifies OoD objects, with anomaly scores visualized from low (blue) to high (red).
Such anomalies include unexpected obstacles (e.g., discarded furniture, stray wildlife, construction equipment), unusual road conditions (e.g., potholes, debris, or oil spills), and atypical weather (e.g., sudden fog, heavy rain, or snowdrifts). As existing systems may misinterpret or fail to detect these deviations, they pose significant safety risks, potentially leading to failures and even catastrophic consequences in autonomous driving.
@@
To bridge this critical gap, we introduce the task of Out-of-Distribution Semantic Occupancy Prediction, aiming to advance semantic scene understanding by integrating 3D occupancy prediction with OoD detection. However, the lack of datasets annotated for both semantic occupancy and OoD objects hinders robust model development. To address this, we propose a Realistic Anomaly Augmentation that injects diverse anomalies into existing datasets, preserving realistic spatial and occlusion patterns. This pipeline enables the creation of two novel datasets, VAA-KITTI and VAA-KITTI-360, comprising $26$ distinct categories of anomalies, where anomaly annotations encompass a diverse range of unforeseen hazards, facilitating comprehensive evaluation of OoD detection in 3D scene understanding. To explore real-world generalization, we further contribute VAA-STU, a real-world anomaly dataset for evaluation purposes.
@@
Building upon our newly proposed datasets, we introduce OccOoD, the first unified framework that seamlessly unifies semantic occupancy prediction with OoD detection. First, we propose Cross-Space Semantic Refinement (CSSR), which refines semantic predictions by integrating complementary voxel and BEV features, where BEV provides enhanced global context to supplement local voxel representations, thereby improving both occupancy prediction robustness and OoD detection effectiveness. Second, we introduce an OoD detection mechanism leveraging entropy and cosine similarity, enabling effective identification of anomalies in dynamic, unstructured scenes. Furthermore, to ensure comprehensive and realistic evaluation, we conduct extensive OoD detection experiments on the newly proposed datasets and validate on the real-world VAA-STU dataset, while evaluating occupancy prediction on public benchmarks. Experimental results show that OccOoD achieves AuROC scores of $65.50\%$ and $47.90\%$ on the proposed datasets, $65.37\%$ on VAA-STU, and mIoU scores of $16.80\%$ and $18.37\%$ on SemanticKITTI [2] and SSCBench-KITTI-360 [27], respectively. These results collectively establish OccOoD as a robust solution for addressing OoD challenges in 3D semantic occupancy prediction, demonstrating generalization to real-world environments.
In a nutshell, our contributions are threefold:
@@
  • We for the first time introduce out-of-distribution detection into the task of semantic occupancy prediction by delivering two new tailored datasets, VAA-KITTI and VAA-KITTI-360, specifically designed to evaluate OoD detection in the context of autonomous driving safety.
@@
  • We propose OccOoD, a unified framework that jointly handles 3D semantic occupancy prediction and OoD detection by fusing front-view and bird’s-eye-view representations, which provides the comprehensive spatial awareness required to robustly identify outliers.
@@
  • Our proposed framework, OccOoD, achieves state-of-the-art OoD detection performance on the established datasets while maintaining competitive accuracy in semantic occupancy prediction.
II Related Work
@@
Semantic Occupancy Prediction. 3D semantic occupancy prediction aims to infer both the geometry and semantics of a scene by inferring the occupancy state of voxels [42]. To advance estimation quality and efficiency, efforts are made in 3D representation [30, 33], context enhancement [47, 24], modality fusion [52, 34], and long-sequence and generative modeling [46, 23]. These efforts have collectively improved the quality and efficiency of semantic occupancy prediction. Recently, several works [19, 50] have introduced more fine-grained feature processing techniques. Meanwhile, uncertainty-aware methods [21, 17] incorporate uncertainty quantification to enhance estimation reliability. However, these methods are limited to predefined categories, requiring retraining to generalize to novel object types and lacking flexibility. While open-vocabulary learning [43, 8, 45, 54] addresses generalization to arbitrary categories, it remains constrained to In-Distribution (ID) scenarios and fails to handle OoD detection. In real-world applications, the categories of anomalous objects are unforeseeable, making such methods less reliable. In contrast, our proposed model extends semantic occupancy prediction by perceiving all OoD objects, enabling efficient and comprehensive scene understanding.
@@
Figure 2: An illustration of the proposed Realistic Anomaly Augmentation. This pipeline is designed to address the challenges of collecting real-world OoD data by synthesizing anomalies that are physically plausible and contextually realistic.
@@
Out-of-Distribution Segmentation. Out-of-distribution segmentation identifies anomalies in specific scenes [10, 7] while maintaining in-distribution segmentation capabilities. Current research primarily focuses on anomaly detection in 2D images of road scenes, with the main methods including per-pixel architectures and mask-transformer-based approaches [41, 38]. The per-pixel architecture methods include uncertainty-based [11, 20], reconstruction-based [29, 15], and outlier exposure [44, 16] methods. While advancing 2D detection, these methods are constrained by pinhole camera inputs, which limits their applicability in 3D scenarios demanding precise spatial perception. Some studies extend anomaly detection to 3D space, yet primarily target industrial [3, 31] or indoor object inspection [6, 25], typically relying on static point clouds for defect localization [4, 51]. Due to their dependence on static geometric consistency and limited responsiveness to dynamic changes, these approaches are unsuitable for unconstrained, evolving environments. In contrast, we integrate out-of-distribution segmentation into vision-based 3D semantic occupancy prediction, enabling effective response to real-time changes and uncertainties, thereby overcoming the limitations of current methods in dynamic settings.
III Out-of-Distribution Datasets
To address the lack of evaluation platforms, we propose a Realistic Anomaly Augmentation to generate abnormal objects in real road scene images, leading to the creation of a suite of datasets: VAA-KITTI, VAA-KITTI-360, and VAA-STU. The former two are generated with realistic spatial and occlusion patterns, while VAA-STU is derived from a real-world 3D point cloud dataset [39] containing natural outliers, with scenes voxelized for occupancy prediction, enabling evaluation under real-world OoD conditions.
III-A Realistic Anomaly Augmentation
@@
Collecting images and point cloud data of OoD objects is challenging due to their rarity in real-world scenes and the high cost of data collection and annotation. To address this, we propose a Realistic Anomaly Augmentation that generates synthetic anomalies under physical and environmental constraints, ensuring plausibility and challenge for robust OoD detection model evaluation. As illustrated in Fig. 2, our pipeline is structured into three key phases:
@@
Phase 1: Anomaly Generation and Localization. In 2D anomaly detection, studies have demonstrated that synthetic anomaly data is highly effective for model training and evaluation [12, 13]. We employ POC [12], a state-of-the-art image generation technique, to synthesize anomaly images and their corresponding mask maps. To better align with real-world application scenarios, we introduce multi-category anomalies into the images, as illustrated in Fig. 3.
@@
Phase 2: Depth Estimation and 3D Transformation. This phase integrates images with accurate depth to enable transformation of 2D anomaly detections into precise 3D spatial localization. To address limited metric depth accuracy, we project synthesized anomalies into the 3D point cloud.
@@
Specifically, we first utilize the large version of Depth Anything V2 [48], to compute pseudo-depth for both original and synthesized images. Subsequently, we randomly sample a predefined number of points, from the point cloud and project them onto the image plane to obtain pixel coordinates $\mathbf{u}_{i}$ and Euclidean distances $d_{i}$ as:
@@
where 0 and $1$ represent the occupancy status of the voxel.
III-B VAA-KITTI, VAA-KITTI-360, and VAA-STU
Figure 3: Distribution map of VAA-KITTI and VAA-KITTI-360, containing 2626 distinct OoD categories grouped into five main types as shown.
We applied the aforementioned synthesis pipeline to SemanticKITTI [2] and SSCBench-KITTI-360 [27], creating two fresh datasets: VAA-KITTI and VAA-KITTI-360. These datasets extend the original labels with outliers, encompassing 26 distinct anomaly categories.
To ensure the plausibility and naturalness, over $30,000$ images produced using the POC algorithm were carefully reviewed by three independent annotators. After multiple rounds of cross-checking, images with poor generation quality, implausible object placements, or non-anomalous scenarios (e.g., birds flying in an urban sky or trash and wheeled bins placed alongside sidewalks) are excluded, yielding a final selection of $500$ images for both the VAA-KITTI and VAA-KITTI-360 test sets in the Out-of-Distribution Semantic Occupancy Prediction task.
In addition, we construct VAA-STU from a real-world 3D LiDAR dataset featuring real anomalous objects with precise instance annotations. It includes two sequences from STU [39], voxelized into a format compatible with occupancy models, preserving real-world environmental complexity for the robust OoD evaluation.
Regional-Level Metrics Design. Referring to 2D image anomaly detection [38], we adopt Area under the Receiver Operating Characteristic Curve (AuROC) as a performance metric. Furthermore, to mitigate the evaluation impact caused by the inaccuracy of existing methods in predicting the absolute positions of instances, and bridge the gap between localization demands and anomaly-awareness goals, we propose a regional-level metric: Regional Area under the Precision-Recall Curve ($\text{AuPRC}_{r}$). The $\text{AuPRC}_{r}$ incorporates tolerance for minor localization inaccuracies while maintaining sensitivity to anomaly detection.
$\text{AuPRC}_{r}$: The Precision-Recall curve evaluates the balance between precision and recall at varying thresholds:
@@ III-B VAA-KITTI, VAA-KITTI-360, and VAA-STU
IV OccOoD: Proposed Framework
IV-A Problem Setting
@@
Out-of-distribution segmentation in 2D pinhole images typically relies on pixel-wise anomaly scores. Building on this concept, we extend out-of-distribution detection to 3D semantic occupancy prediction at the voxel level. Given an input image, the semantic occupancy prediction process yields a volumetric representation $V\in\mathbb{R}^{C\times H\times W\times Z}$, where $C$, $H$, $W$, and $Z$ correspond to the channels, height, width, and depth of the voxel space, respectively. The voxel-level classification scores $Q(V)\in\mathbb{R}^{K\times H\times W\times Z}$ are then obtained, and $K$ denotes the number of classes for the semantic prediction. Ultimately, to assess anomalies at the voxel level, an anomaly score $A(V)\in\mathbb{R}^{H\times W\times Z}$ is computed as Eq. (6):
$A(V)=f(Q(V)).$ (6)
IV-B Overall Architecture